# PFGuard: A Generative Framework with Privacy and Fairness Safeguards

### v1

@@ Abstract
Generative models must ensure both privacy and fairness for Trustworthy AI. While these goals have been pursued separately, recent studies propose to combine existing privacy and fairness techniques to achieve both goals. However, naïvely combining these techniques can be insufficient due to privacy-fairness conflicts, where a sample in a minority group may be amplified for fairness, only to be suppressed for privacy. We demonstrate how these conflicts lead to adverse effects, such as privacy violations and unexpected fairness-utility tradeoffs. To mitigate these risks, we propose PFGuard, a generative framework with privacy and fairness safeguards, which simultaneously addresses privacy, fairness, and utility. By using an ensemble of multiple teacher models, PFGuard balances privacy-fairness conflicts between fair and private training stages and achieves high utility based on ensemble learning. Extensive experiments show that PFGuard successfully generates synthetic data on high-dimensional data while providing both fairness convergence and strict DP guarantees – the first of its kind to our knowledge.
@@ 1 Introduction
Recently, generative models have shown remarkable performance in various applications including vision (Wang et al., 2021b) and language tasks (Brown et al., 2020) – while also raising significant ethical concerns. In particular, privacy and fairness concerns have emerged due to generative models mimicking their training data. On the privacy side, specific training data can be memorized, allowing the reconstruction of sensitive information like facial images (Hilprecht et al., 2019; Sun et al., 2021). On the fairness side, any bias in the training data can be learned, resulting in biased synthetic data and unfair downstream performances across demographic groups (Zhao et al., 2018; Tan et al., 2020).
@@
Figure 1: Privacy-fairness conflict. Privacy techniques prefer the left-hand scenario to prevent privacy risk of a certain data sample, while fairness techniques prefer the right-hand scenario to balance learning w.r.t. groups. Related empirical results are shown in Table 5.2.
@@
Figure 2: Overview of PFGuard. PFGuard integrates fairness and privacy in generative models through a two-stage process. In the fair training stage (blue), we train fair teacher models by sampling balanced mini-batches from biased data subsets. In the private training stage (red), we aggregate the teacher outputs to supervise the generator (e.g., whether the generated image given input noise $Z$ is realistic or unrealistic) with DP noise to ensure privacy. With these training stages to achieve fairness and privacy, PFGuard also achieves high utility based on ensemble learning of teacher models, resulting in high-quality, unbiased, and private synthetic data.
Therefore, we design a generative framework that simultaneously addresses fairness and privacy while achieving utility for high-dimensional synthetic data generation. To this end, we propose PFGuard: a generative framework with Privacy and Fairness Safeguards. As illustrated in Fig. 2, the key component is an ensemble of intermediate teacher models, which balances privacy-fairness conflicts between fair training and private training stages. In the fair training stage, we design a new sampling technique to train fair teachers, which provides a theoretical convergence guarantee to the fair generator. In the private training stage, we employ the Private Teacher Ensemble Learning (PTEL) approach (Papernot et al., 2016; 2018), which aggregates each teacher’s knowledge with random DP noise (e.g., noisy voting), to privatize the knowledge transfer to the generator. As a result, PFGuard provides a unified solution to train both fair and private generative models by transferring the teachers’ fair knowledge in a privacy-preserving manner.
@@
Moreover, PFGuard is compatible with a wide range of existing private generative models and preserves their utility. PTEL is widely adopted in private generative models as it provides prominent privacy-utility tradeoff (Jordon et al., 2018; Chen et al., 2020; Long et al., 2021; Wang et al., 2021a). PFGuard can extend any of these models with a fair training stage as shown in Fig. 2, which requires a simple modification in the minibatch sampling process. Since additional fair sampling does not require additional training complexity compared to say adding a loss term for fairness, PFGuard preserves the privacy-utility tradeoff of PTEL as well while improving fairness. We also provide guidelines to control the fairness-privacy-utility tradeoff – see more details in Sec. 4.
Experiments show that PFGuard successfully generates high-dimensional synthetic data while ensuring both privacy and fairness; to our knowledge, PFGuard is the first framework that works on high-dimensional data including images. Our results also reveal two key findings: (1) existing private generative models can produce highly-biased synthetic data in real-world scenarios even with simple bias settings, and (2) a naïve combination of individual techniques may fail to achieve either privacy or fairness even with simple datasets – highlighting PFGuard’s effectiveness and the need for a better integration of fair and private generative models.
@@ 2 Preliminaries
We focus on Generative Adversarial Networks (Goodfellow et al., 2014), which are widely-used generative models that leverage adversarial training of two networks to generate realistic synthetic data: 1) a generator that learns the underlying training data distribution and generates new samples and 2) a discriminator that distinguishes between real and generated data. The discriminator can be considered as the teacher model of the generator, as the generator does not have access to the real data and only learns from the discriminator via the GAN loss function.
@@
We use Differential Privacy (DP) (Dwork et al., 2014) to privatize generative models, a gold standard privacy notion that provides quantitative privacy analysis of an algorithm. DP measures how much an adversary can learn about a single sample by differentiating two outputs of an algorithm $\mathcal{M}$. This privacy guarantee is quantified with the parameters $(\varepsilon,\delta$) as follows:
@@
($(\epsilon,\delta)$-Differential Privacy (Dwork et al., 2006)) A randomized mechanism $\mathcal{M}:\mathcal{D}\rightarrow\mathcal{R}$ with domain $\mathcal{D}$ and range $\mathcal{R}$ satisfies $(\epsilon,\delta)$-differential privacy if for any two adjacent inputs $D,D^{\prime}$, which differ by a single sample, and for any subset of outputs $O\subseteq\mathcal{R}$, the following holds:
$\Pr(\mathcal{M}(D)\in O)\leq e^{\epsilon}\Pr(\mathcal{M}(D^{\prime})\in O)+\delta,$
where $\epsilon$ is the upper bound of privacy loss, and $\delta$ is the probability of breaching DP constraints.
We can enforce DP in an algorithm in two steps (Dwork et al., 2014). First, we bound its “sensitivity” (Def. 2.2), which captures the maximum influence of an individual sample. Second, we add random noise that is proportional to the sensitivity. A common way to ensure DP is to use a Gaussian mechanism (Dwork et al., 2014) (Thm. 2.1), which utilizes Gaussian random noise with a scale proportional to $l_{2}$-sensitivity. Two datasets $D$, $D^{\prime}$ are adjacent if they differ by a single sample.
@@
(Sensitivity (Dwork et al., 2014)) The $l_{p}$-sensitivity for a $d$-dimensional function $f:\mathcal{D}{\rightarrow}\mathbb{R}^{d}$ is defined as $\Delta_{f}^{p}=\underset{D,D^{\prime}}{\max}\left\|f(D)-f(D^{\prime})\right\|_{p}$ over all adjacent datasets $D,D^{\prime}$.
@@
(Gaussian mechanism (Dwork et al., 2014; Mironov, 2017)) Let $f:\mathcal{X}\rightarrow\mathbb{R}^{d}$ be an arbitrary $d$-dimensional function with $l_{2}$-sensitivity $\Delta_{f}^{2}$. The Gaussian mechanism $\mathcal{M}_{\sigma}$, parameterized by $\sigma$, adds Gaussian noise into the output, i.e., $\mathcal{M}_{\sigma}({\bm{x}})=f({\bm{x}})+\mathcal{N}(0,\sigma^{2}\bm{I})$, and satisfies $(\varepsilon,\delta)$-DP for $\sigma\geq\sqrt{2\ln(1.25/\delta)}\Delta_{f}^{2}/\varepsilon$.
@@
3 Challenges of Satisfying Both Privacy and Fairness
In this section, we examine the practical challenges of integrating privacy-only and fairness-only techniques to train both private and fair generative models. Based on Fig. 1’s intuition on how privacy and fairness conflict, we analyze how existing approaches for DP generative models and fair generative models can technically conflict with each other.
Adding Fairness Can Worsen Privacy
Ensuring fairness in DP generative models can significantly increase sensitivity (Def. 2.2), leading to invalid DP guarantees. Sensitivity, which measures a data sample’s maximum impact on an algorithm, is crucial in DP generative models because the noise amount required for DP guarantees is often proportional to this sensitivity value. However, integrating fairness techniques in DP generative models can invalidate their sensitivity analyses by adjusting model outputs for fairness purposes. Examples include amplifying the impact of certain data samples to balance model training across groups (Choi et al., 2020) and directly feeding data attributes such as class labels or sensitive attributes (e.g., race, gender) to a generator for more balanced synthetic data (Xu et al., 2018; Sattigeri et al., 2019; Yu et al., 2020), which can cause large fluctuations in the generator output with any variation in these data attributes. As a result, fairness techniques can end up increasing sensitivity and require more noise to maintain the same privacy level, compromising the original DP guarantees unless modifying DP techniques to add more noise. However, this modification is also not straightforward as assessing the increased sensitivity by fairness techniques can be challenging (Tran et al., 2021b).
Adding Privacy Can Worsen the Fairness-Utility Tradeoff
Another direction is to ensure privacy in fair generative models, but configuring an appropriate privacy bound can be challenging, leading to unexpected fairness-utility tradeoffs. We illustrate below using standard DP and fairness techniques: DP-SGD (Abadi et al., 2016) and reweighting (Choi et al., 2020). Let $\bm{g}({\mathbf{x}})$ denote the gradient of the data sample ${\mathbf{x}}$, and $p_{\text{bal}}$ and $p_{\text{bias}}$ denote balanced and biased data distributions, respectively.
@@ 3 Challenges of Satisfying Both Privacy and Fairness
We can extend reweighting for fairness to also satisfy DP using DP-SGD, but finding the clipping threshold $C$ that balances fairness and utility can be challenging. Note that if we perform DP-SGD and then reweighting, privacy breach may occur by amplifying sample gradients beyond $C$, invalidating the sensitivity derived from gradient clipping. We thus consider performing reweighting and then DP-SGD, which at least guarantees DP for reweighting-based fair generative models. However, the clipping now undoes the fairness adjustments where reweighted gradients $\bm{g}({\mathbf{x}})\cdot h({\mathbf{x}})$ are clipped to $\bm{g}({\mathbf{x}})\cdot h({\mathbf{x}})/\max(1,||\bm{g}({\mathbf{x}})\cdot h({\mathbf{x}})||_{2}/C)$, and Eq. 1 does not hold if $C\leq\bm{g}({\mathbf{x}})$. Here one solution is to use a larger $C$ such that $C\geq\bm{g}({\mathbf{x}})$. However, increasing $C$ also increases the noise required for DP, which reduces utility (Fig. 3). As a result, selecting a $C$ that balances fairness and utility may necessitate extensive hyperparameter tuning (Bu et al., 2024), complicating the systematic integration of fairness into DP generative models.
Overall, we show that a naïve combination of existing fairness-only and privacy-only techniques can be insufficient to achieve both objectives. While we have not exhaustively covered all possible combinations, one can see how privacy breaches and unexpected fairness-utility tradeoffs can easily occur without a careful design. To avoid these downsides, we emphasize the need for a unified design that integrates both privacy and fairness in generative models.
Remark 1.
Our study is the first to reveal that fairness and privacy techniques can counteract each other. We also demonstrate how both can be compromised – see empirical results in Sec. 5.
@@ 4 Framework
We now propose PFGuard, the first generative framework that simultaneously achieves statistical fairness and DP on high-dimensional data, such as images. As shown in Fig. 2, PFGuard balances privacy-fairness conflicts between fair and private training stages using an ensemble of teacher models as a key component. In Sec. 4.1, we first explain the fair training stage, which trains a fair teacher ensemble. In Sec. 4.2, we then explain the private training stage, which transfers the knowledge of this teacher ensemble to the generator with DP guarantees – ultimately training a generator that is both fair and private. In Sec. 4.3, we lastly discuss how PFGuard’s integrated design offers advantages in terms of fairness, utility, and privacy compared to the naïve approaches discussed in Sec. 3.
@@ 4.1 Fair Training with Balanced Minibatch Sampling
We ensure fairness in the teachers by balancing the minibatches used for training. Here we assume a general training setup of stochastic gradient descent, where we iteratively pick a subset $B$ of the training data (i.e., minibatches) to update model parameters more efficiently. Since generative models then only learn the underlying data distribution through $B$, using balanced minibatches $B\sim p_{\text{bal}}$ will result in modeling $p_{\text{bal}}$ even if we have biased training data $D\sim p_{\text{bias}}$. While another approach is to debias the training data itself by acquiring more minority data, this approach is costly and often infeasible for private domains with limited publicly-available data (Jordon et al., 2018).
@@
To develop a fair minibatch sampling technique with a convergence guarantee, we leverage Sampling-Importance Resampling (SIR) (Rubin, 1988; Smith & Gelfand, 1992) as the theoretical foundation. SIR is a statistical method for drawing random samples from a target distribution $\pi(x)$ by using a proposal distribution $\psi(x)$. SIR proceeds in two steps: 1) we draw a set of $n$ independent random samples $S_{1}=\{x_{i}\}_{i=1}^{n}$ from $\psi(x)$ and 2) we resample a smaller set of $m$ independent random samples $S_{2}=\{x_{i}\}_{i=1}^{m}$ from $S_{1}$. Here, the resampling probability $w(x_{i})$ is set proportional to $h(x_{i})$, where $h(x_{i})=\pi(x_{i})/\psi(x_{i})$. Then, the resulting samples in $S_{2}$ are approximately distributed according to $\pi(x)$ as follows:
@@
We now present our sampling technique, which guarantees $B\sim p_{\text{bal}}$ based on SIR. We first make the following reasonable assumptions: 1) each data sample has a uniquely defined sensitive attribute ${\mathbf{s}}\in\mathcal{S}$ (e.g., race); 2) $p_{\text{bal}}$ is uniformly distributed over ${\mathbf{s}}$; 3) following Choi et al. (2020), the same relevant input features are shared for each group ${\mathbf{s}}$ between the balanced and biased datasets (e.g., $p_{\text{bal}}({\mathbf{x}}|{\mathbf{s}}{=}s)=p_{\text{bias}}({\mathbf{x}}|{\mathbf{s}}{=}s)$), and similarly between the training dataset $D$ and any subset $D_{i}$ (e.g., $p_{D}({\mathbf{x}}|{\mathbf{s}}{=}s)=p_{D_{i}}({\mathbf{x}}|{\mathbf{s}}{=}s)$). We now outline the technique step-by-step below.
@@
    We set the target distribution to $p_{\text{bal}}({\mathbf{x}})$ and the proposal distribution to $p_{\text{bias}}({\mathbf{x}})$, as our goal is to sample a balanced minibatch $B\sim p_{\text{bal}}$ from the biased training dataset $D\sim p_{\text{bias}}$.
@@
    We divide $D$ into $n_{T}$ disjoint subsets $\{D_{i}\}^{n_{T}}_{i=1}$ to train each teacher model $T_{i}$, such that each $D_{i}$ retains the same value distribution of ${\mathbf{s}}$ as $D$ using the ${\mathbf{s}}$ labels (i.e., $p_{D_{i}}({\mathbf{s}}=s)=p_{D}({\mathbf{s}}=s)$). Then, we can derive $D_{i}\sim p_{\text{bias}}$ using assumption 3) above as follows:
    $p_{D_{i}}({\mathbf{x}})=\Sigma_{s}p_{D_{i}}({\mathbf{x}}|{\mathbf{s}}{=}s)p_{D_{i}}({\mathbf{s}}{=}s)=\Sigma_{s}p_{D}({\mathbf{x}}|{\mathbf{s}}{=}s)p_{D}({\mathbf{s}}{=}s)=p_{\text{bias}}({\mathbf{x}})$ (4)

### v2

@@ Abstract
Generative models must ensure both privacy and fairness for Trustworthy AI. While these goals have been pursued separately, recent studies propose to combine existing privacy and fairness techniques to achieve both goals. However, naïvely combining these techniques can be insufficient due to privacy-fairness conflicts, where a sample in a minority group may be represented in ways that support fairness, only to be suppressed for privacy. We demonstrate how these conflicts lead to adverse effects, such as privacy violations and unexpected fairness-utility tradeoffs. To mitigate these risks, we propose PFGuard, a generative framework with privacy and fairness safeguards, which simultaneously addresses privacy, fairness, and utility. By using an ensemble of multiple teacher models, PFGuard balances privacy-fairness conflicts between fair and private training stages and achieves high utility based on ensemble learning. Extensive experiments show that PFGuard successfully generates synthetic data on high-dimensional data while providing both DP guarantees and convergence in fair generative modeling.
@@ 1 Introduction
Recently, generative models have shown remarkable performance in various applications including vision (Wang et al., 2021b) and language tasks (Brown et al., 2020), while also raising significant ethical concerns. In particular, privacy and fairness concerns have emerged as generative models aim to mimic their training data. On the privacy side, specific training data can be memorized, allowing the leakage of personal sensitive information (Hilprecht et al., 2019; Sun et al., 2021). On the fairness side, any bias in the training data can be learned, resulting in biased synthetic data and unfair downstream performances across demographic groups (Zhao et al., 2018; Tan et al., 2020).
@@
Figure 1: Privacy-fairness conflict. Privacy techniques prefer the left-hand scenario to prevent privacy risk of a certain data sample, while fairness techniques prefer the right-hand scenario to balance learning w.r.t. groups.
@@
Figure 2: Overview of PFGuard. PFGuard integrates fairness and privacy into generative models through a two-stage process. In the fair training stage (blue), we train fair teacher models by sampling balanced mini-batches from biased data partitions. In the private training stage (red), we aggregate the teacher outputs with random noise to ensure Differential Privacy (DP). After training, only the trained DP generator is publicly released, safeguarding the privacy of all other components (e.g., the teacher ensemble). Through these training stages, PFGuard not only ensures fairness and privacy but also achieves high utility by leveraging ensemble learning of teacher models – resulting in unbiased, private, and high-quality synthetic data. More details on the framework are presented in Sec. 4.
Therefore, we design a generative framework that simultaneously addresses fairness and privacy while achieving utility for high-dimensional synthetic data generation. To this end, we propose PFGuard: a generative framework with Privacy and Fairness Safeguards. As illustrated in Fig. 2, the key component is an ensemble of intermediate teacher models, which balances privacy-fairness conflicts between fair training and private training stages. In the fair training stage, we design a new sampling technique to train fair teachers, which provides a theoretical convergence guarantee to the fair generative modeling. In the private training stage, we employ the Private Teacher Ensemble Learning (PTEL) approach (Papernot et al., 2016; 2018), which aggregates each teacher’s knowledge with random DP noise (e.g., noisy voting), to privatize the knowledge transfer to the generator. As a result, PFGuard provides a unified solution to train both fair and private generative models by transferring the teachers’ fair knowledge in a privacy-preserving manner.
@@
Moreover, PFGuard is compatible with a wide range of existing private generative models and preserves their utility. PTEL is widely adopted in private generative models as it provides prominent privacy-utility tradeoff (Jordon et al., 2018; Chen et al., 2020; Long et al., 2021; Wang et al., 2021a). PFGuard can extend any of these models with a fair training stage as shown in Fig. 2, which requires a simple modification in the minibatch sampling process. Since additional fair sampling can be advantageous in maintaining optimization complexity compared to say adding a loss term for fairness, PFGuard preserves the privacy-utility tradeoff of PTEL as well while improving fairness. We also provide guidelines to control the fairness-privacy-utility tradeoff – see more details in Sec. 4.
Experiments in Sec. 5 show that PFGuard successfully generates high-dimensional synthetic data while ensuring both privacy and fairness; to our knowledge, PFGuard is the first framework that works on high-dimensional data including images. Our results also reveal two key findings: (1) existing private generative models can produce highly-biased synthetic data in real-world scenarios even with simple bias settings, and (2) a naïve combination of individual techniques may fail to achieve either privacy or fairness even with simple datasets – highlighting PFGuard’s effectiveness and the need for a better integration of privacy and fairness in generative models.
@@ 2 Preliminaries
We focus on Generative Adversarial Networks (GAN) (Goodfellow et al., 2014), which are widely-used generative models that leverage adversarial training of two networks to generate realistic synthetic data: 1) a generator that learns the underlying training data distribution and generates new samples and 2) a discriminator that distinguishes between real and generated data. The discriminator can be considered as the teacher model of the generator, as the generator does not have access to the real data and only learns from the discriminator via the GAN loss function.
@@
To privatize generative models, we use Differential Privacy (DP) (Dwork et al., 2014), a gold standard privacy framework that enables quantified privacy analyses. DP measures how much an adversary can infer about one data sample based on differences in two outputs of an algorithm $\mathcal{M}$, using two adjacent datasets that differ by one sample as defined below.
@@
($(\varepsilon,\delta)$-Differential Privacy (Dwork et al., 2006a)) A randomized mechanism $\mathcal{M}:\mathcal{D}\rightarrow\mathcal{R}$ with range $\mathcal{R}$ satisfies $(\varepsilon,\delta)$-differential privacy if for any two adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}$ and for any subset of outputs $\mathcal{O}\subseteq\mathcal{R}$, the following holds:
$\Pr(\mathcal{M}(\mathcal{D})\in\mathcal{O}\leq e^{\varepsilon}\Pr(\mathcal{M}(\mathcal{D}^{\prime})\in\mathcal{O})+\delta,$
where $\varepsilon$ is the upper bound of privacy loss, and $\delta$ is the probability of breaching DP constraints.
We can enforce DP in an algorithm in two steps (Dwork et al., 2014). Given a target algorithm $f$ to enforce DP and a dataset $\mathcal{D}$, we first bound sensitivity (Def. 2.2), which captures the maximum influence of a single data sample on the output of $f$. We then add random noise with a scale proportional to the sensitivity value. A common way to add noise is to utilizing a Gaussian mechanism (Dwork et al., 2014) (Thm. 2.1), which uses Gaussian random noise with a scale proportional to $l_{2}$-sensitivity.
@@
(Sensitivity (Dwork et al., 2014)) The $l_{p}$-sensitivity for a $d$-dimensional function $f:X\rightarrow\mathbb{R}^{d}$ is defined as $\Delta_{f}^{p}=\underset{\mathcal{D},\mathcal{D}^{\prime}}{\max}\left\|f(\mathcal{D})-f(\mathcal{D}^{\prime})\right\|_{p}$ over all adjacent datasets $\mathcal{D},\mathcal{D}^{\prime}$.
@@
(Gaussian mechanism (Dwork et al., 2014; Mironov, 2017)) Let $f:X\rightarrow\mathbb{R}^{d}$ be an arbitrary $d$-dimensional function with $l_{2}$-sensitivity $\Delta_{f}^{2}$. The Gaussian mechanism $\mathcal{M}_{\sigma}$, parameterized by $\sigma$, adds Gaussian noise into the output, i.e., $\mathcal{M}_{\sigma}({\bm{x}})=f({\bm{x}})+\mathcal{N}(0,\sigma^{2}\bm{I})$, and satisfies $(\varepsilon,\delta)$-DP for $\sigma\geq\sqrt{2\ln(1.25/\delta)}\Delta_{f}^{2}/\varepsilon$.
@@
3 Privacy-Fairness Conflicts in Existing Techniques
In this section, we examine the practical challenges of integrating privacy-only and fairness-only techniques to train both private and fair generative models. Based on Fig. 1’s intuition on how privacy and fairness conflict, we analyze how existing approaches for DP generative models and fair generative models can technically conflict with each other using standard DP and fairness techniques: DP-SGD (Abadi et al., 2016) and reweighting (Choi et al., 2020). Let $\bm{g}({\mathbf{x}})$ denote the gradient of the data sample ${\mathbf{x}}$, and $p_{\text{bal}}$ and $p_{\text{bias}}$ denote balanced and biased data distributions, respectively.
@@ 3 Privacy-Fairness Conflicts in Existing Techniques
Naïvely Adding Fairness Can Worsen Privacy
Ensuring fairness in DP generative models can significantly increase sensitivity (Def. 2.2), leading to invalid DP guarantees. Sensitivity, which measures a data sample’s maximum impact on an algorithm, is crucial in DP generative models because the noise amount required for DP guarantees is often determined by this sensitivity value. However, integrating fairness techniques in DP generative models can invalidate their sensitivity analyses by adjusting model outputs for fairness purposes. One example is the aforemetioned reweighting technique. To balance model training across groups, reweighting technique amplifies the impact of certain data samples (e.g., minority data samples), particularly by amplifying their gradients. However, performing reweighting on top of DP-SGD can incur privacy breaches, as gradient amplified beyond $C$ can cancel the gradient clipping in DP-SGD, invalidating the sensitivity $C$ provided from DP-SGD. Other examples include directly feeding data attributes such as class labels or sensitive attributes (e.g., race, gender) to a generator for more balanced synthetic data (Xu et al., 2018; Sattigeri et al., 2019; Yu et al., 2020), which can cause large fluctuations in the generator output and similarly end up increasing sensitivity. This increased sensitivity by fairness techniques requires more noise to maintain the same privacy level, compromising the original DP guarantees unless modifying DP techniques to add more noise. However, this modification is also not straightforward as assessing the increased sensitivity by fairness techniques can be challenging (Tran et al., 2021b).
Naïvely Adding Privacy Can Worsen the Fairness-Utility Tradeoff
Another direction is to ensure privacy in fair generative models, but configuring an appropriate privacy bound can be challenging, which can lead to unexpected and unstable fairness-utility tradeoffs.
@@
In contrast to the previous demonstration, we can ensure privacy in reweighting-based fair generative models by replacing SGD to DP-SGD, guaranteeing a fixed sensitivity value $C$ via gradient clipping; however, finding $C$ that balances fairness and utility can be challenging. Gradient clipping now undoes the fairness adjustments, as reweighted gradients $\bm{g}({\mathbf{x}}){\cdot}h({\mathbf{x}})$ are clipped to $\bm{g}({\mathbf{x}}){\cdot}h({\mathbf{x}})/\max(1,||\bm{g}({\mathbf{x}}){\cdot}h({\mathbf{x}})||_{2}/C)$, and Eq. 1 does not hold if $C\leq\bm{g}({\mathbf{x}}){\cdot}h({\mathbf{x}})$. Here, one solution is to use a larger $C$ such that $C\geq\bm{g}({\mathbf{x}}){\cdot}h({\mathbf{x}})$. However, increasing $C$ also increases the noise required for DP, which reduces utility as illustrated in Fig. 3. Therefore, selecting a $C$ that balances fairness and utility may necessitate extensive hyperparameter tuning (Bu et al., 2024), complicating the systematic integration of DP into fair generative models.
Overall, we show that a naïve combination of existing fairness-only and privacy-only techniques can be insufficient to achieve both objectives. While we have not exhaustively covered all possible combinations, one can see how privacy breaches and unexpected fairness-utility tradeoffs can easily occur without a careful design. To avoid these downsides in naïve combinations, we emphasize the need for a unified design that integrates both privacy and fairness into generative models.
@@ 4 Framework
We now propose PFGuard, the first generative framework that simultaneously achieves statistical fairness and DP on high-dimensional data, such as images. As shown in Fig. 2, PFGuard balances privacy-fairness conflicts between fair and private training stages using an ensemble of teacher models as a key component. In Sec. 4.1, we first explain the fair training stage, which trains a fair teacher ensemble. In Sec. 4.2, we then explain the private training stage, which transfers the knowledge of this teacher ensemble to the generator with DP guarantees and ultimately trains a generator that is both fair and private. In Sec. 4.3, we lastly discuss how PFGuard’s integrated design offers advantages in terms of fairness, utility, and privacy compared to the naïve approaches discussed in Sec. 3.
@@ 4.1 Fair Training with Balanced Minibatch Sampling
We ensure fairness in the teachers by balancing the minibatches used for training. Here we assume a general training setup of stochastic gradient descent, where we iteratively pick a subset $\mathcal{B}$ of the training data (i.e., minibatches) to update model parameters more efficiently. Since generative models then only learn the underlying data distribution through $\mathcal{B}$, using balanced mini-batches $\mathcal{B}\sim p_{\text{bal}}$ will result in modeling $p_{\text{bal}}$ even if we have biased training data $\mathcal{D}\sim p_{\text{bias}}$. While another approach is to debias the training data itself by acquiring more minority data, this approach is costly and often infeasible for private domains with limited publicly-available data (Jordon et al., 2018).
@@
To develop a fair minibatch sampling technique with a convergence guarantee, we leverage Sampling-Importance Resampling (SIR) (Rubin, 1988; Smith & Gelfand, 1992) as the theoretical foundation. SIR is a statistical method for drawing random samples from a target distribution $\pi(x)$ by using a proposal distribution $\psi(x)$. SIR proceeds in two steps: 1) we draw a set of $n$ independent random samples $\mathcal{R}_{1}{=}\{x_{i}\}_{i=1}^{n}$ from $\psi(x)$ and 2) we resample a smaller set of $m$ independent random samples $\mathcal{R}_{2}{=}\{x_{i}\}_{i=1}^{m}$ from $\mathcal{R}_{1}$. Here, the resampling probability $w(x_{i})$ is set proportional to $h(x_{i}){=}\pi(x_{i})/\psi(x_{i})$, which is the likelihood ratio of a sample $x_{i}$ in $\pi(x)$ and $\psi(x)$. Then, the resulting samples in $\mathcal{R}_{2}$ are approximately distributed according to $\pi(x)$ as follows:
@@
We now present our sampling technique, which samples $\mathcal{B}\sim p_{\text{bal}}$ based on SIR. We first make the following reasonable assumptions: 1) each data sample has a uniquely defined sensitive attribute ${\mathbf{s}}\in\mathcal{S}$ (e.g., race); 2) target $p_{\text{bal}}$ is a uniform distribution over ${\mathbf{s}}$; 3) following Choi et al. (2020), the same relevant input features are shared for each group ${\mathbf{s}}$ between the balanced and biased datasets (e.g., $p_{\text{bal}}({\mathbf{x}}|{\mathbf{s}}{=}s)=p_{\text{bias}}({\mathbf{x}}|{\mathbf{s}}{=}s)$), and similarly between the training dataset $\mathcal{D}$ and any subset $\mathcal{D}_{i}$ (e.g., $p_{\mathcal{D}}({\mathbf{x}}|{\mathbf{s}}{=}s)=p_{\mathcal{D}_{i}}({\mathbf{x}}|{\mathbf{s}}{=}s)$). We now outline the technique step-by-step below.
@@
    We set the target distribution to $p_{\text{bal}}({\mathbf{x}})$ and the proposal distribution to $p_{\text{bias}}({\mathbf{x}})$, as our goal is to sample a balanced minibatch $\mathcal{B}\sim p_{\text{bal}}$ from the biased training dataset $\mathcal{D}\sim p_{\text{bias}}$.
@@
    We divide $\mathcal{D}$ into $n_{T}$ disjoint subsets $\{\mathcal{D}_{i}\}^{n_{T}}_{i=1}$ to train teacher models $\{T_{i}\}^{n_{T}}_{i=1}$, such that each $\mathcal{D}_{i}$ is used for $T_{i}$ and retains the same distribution of ${\mathbf{s}}$ as $\mathcal{D}$ using the ${\mathbf{s}}$ labels (i.e., $p_{\mathcal{D}_{i}}({\mathbf{s}}{=}s)=p_{\mathcal{D}}({\mathbf{s}}{=}s)$). Then, we can derive $\mathcal{D}_{i}{\sim}p_{\text{bias}}$ using assumption 3) above as follows:
    $p_{\mathcal{D}_{i}}({\mathbf{x}})=\Sigma_{s}p_{\mathcal{D}_{i}}({\mathbf{x}}|{\mathbf{s}}{=}s)p_{\mathcal{D}_{i}}({\mathbf{s}}{=}s)=\Sigma_{s}p_{\mathcal{D}}({\mathbf{x}}|{\mathbf{s}}{=}s)p_{\mathcal{D}}({\mathbf{s}}{=}s)=p_{\text{bias}}({\mathbf{x}})$ (4)